OpenStudy (anonymous):

If you put $1,000 in a savings account that pays 5% interest compounded continuously, how much money will you have in your account in 10 years? Assume you make no additional deposits or withdrawals. (Points : 2) $1,628.89 $1,051.27 $1,648.72 $5,525.85

OpenStudy (wolf1728):

First calculate 5% compounded continuously: annual rate = (e^r) -1 annual rate = (2.718281828^.05) -1 annual rate = 1.0512710964 -1 annual rate = .0512710964 or 5.12710964% total = $1,000 * (1.0512710964)^n total = $1,000 * 1.6487212706 total = 1,648.72 (which is the third choice down)